OpenStudy (anonymous):

What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth.

OpenStudy (anonymous):

10.19803

OpenStudy (anonymous):

change in x is 10, change in y is 2 a^2 + b^2 = c^2 10^2 + 2^2 = c^2 100 + 4 = c^2 sqrt(c) = 104 c = 10.198039

OpenStudy (anonymous):

oops, thats sqrt(104) = c